# Pressure and strain effects on the ab initio electronic structure of La3Ni2O7

### v1

@@ Abstract
The recent discovery of superconductivity in La3Ni2O7 at a critical temperature above 80Β K points to a non-conventional pairing mechanism in nickelates as in cuprates, possibly due to electronic correlations. We have calculated from first principles the electronic structure of La3Ni2O7 under the effect of pressure and epitaxial strain including correlations by the approximation to the many-body self-energy. We find that the Fermi surface is composed of a characteristic cuprate-shape sheet plus a nickelate-specific cylinder , both from Ni orbitals, with a non-negligible drop in the quasiparticle weight and an effective 1D character. This topology results from a delicate balance between the Ni-3 hole pocket , which is suppressed by correlations, and an emerging La-5 electron pocket induced by both correlation and pressure/strain effects and whose role at low energy has been neglected so far. Unlike cuprates, the electronic structure of La3Ni2O7 is already correctly described from ab initio and in agreement with the experiment without the need to introduce Hubbard adjustable parameters or to invoke a strongly correlated physics.
@@ I Introduction
The recent discovery of superconductivity (SC) in Ruddlesden-Popper (RP) nickelates [1, 2, 3] has introduced a new direction in the field of high-temperature superconductivity. Notably, superconductivity has been observed in bulk samples of the bilayer La3Ni2O7 under high pressure [1], with the superconducting phase forming a right-triangular region spanning 14 to 80 GPa, and reaching a maximum onset critical temperature of approximately 80Β K at 18 GPa [4]. More recently, superconductivity has also been reported at Β K in La3Ni2O7 thin films under epitaxial compressive strain at ambient pressure [5, 6]. While the superconducting mechanism is still debated in infinite-layer LnNiO2 nickelates [7, 8, 9, 10, 11], where critical temperatures do not overstep 30Β K [12, 13], the high critical temperature observed in La3Ni2O7 is difficult to reconcile with a BCS electron-phonon pairing mechanism [14, 15], thus pointing to unconventional superconductivity like in cuprates. Indeed, cuprates and nickelates share many analogies: they are both organized in a layered crystal structure with CuO2 or NiO2 planes; and present the same βcuprate-shapeβ Fermi surface sheet of main 3 character, although nickelates bilayer and trilayer show a strong Ni-3 hybridization in the antinodal direction. However, there are also important differences: the constant presence of an anti-ferromagnetic insulating parent phase in cuprates has no analogous in nickelates. For what concerns the electronic structure, the fact that in nickelates the Ni-3 states are closer to the Fermi level than in cuprates, gives rise to at least another Fermi sheet and possibly to a multiband character superconductivity. Therefore, although both unconventional, the pairing mechanism could be possibly different in cuprates and nickelates.
@@
In this work we study from ab initio the electronic structure of La3Ni2O7 including correlations at the level of the approximation to the self-energy [59, 60, 61, 62]. In contrast to the Kohn-Sham eigenvalues used in density-functional theory (DFT), which are the energies of a fictitious non-interacting system, quasiparticle energies have a direct physical interpretation as the excitation energies measured in ARPES. The approximation takes into account electronic exchange and correlations without introducing any adjustable parameter, neither in the DFT functional (e.g.Β an hybrid mixing parameter) nor in the Hamiltonian (e.g.Β a Hubbard interaction term). The weight of correlation effects and their importance in this system is revealed by a comparison between the electronic structure and the DFT band plot.
We also investigate the electronic structure evolution of bulk La3Ni2O7 with respect to pressure, as well as in presence of in-plane constraints that simulate a thin film grown on a substrate. Our purpose is to highlight aspects which might be relevant to explain superconductivity that in this system is induced by pressure without doping. The focus is at the Fermi level but also at higher energy features which are less affected by the estimated 0.1Β eV best absolute accuracy, and which might be relevant in the competition between different pairing symmetries according to some proposed superconducting mechanisms [24]. Our work describes some of the electronic structure features expected to be relevant by theoretical models so far proposed [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48], and provides also indications for future works.
The paper is organized as follows: after introducing our methods in Sec.Β II, we will first compare our calculations with previous ab initio works done at 29.5Β GPa with the crystal structure used by almost all DFT as well as two calculations [45, 15] (Sec.Β III.1). We will then show the evolution with pressure of the electronic structure for the bulk crystal in Sec.Β III.2, and finally introduce in Sec.Β III.3 our results for the structure with epitaxial strain of 1.8%. Our results will be compared with the available experimental data from photoemission (ARPES) experiments.
@@ II.1 Numerical methods
Our methodology relies as a first step on density-functional theory (DFT) calculations in the PBE approximation [63], using the plane waves code Abinit [64]. We use norm-conserving pseudo-potentials from the PseudoDojo table library [65]. The La pseudo-potential takes 4 electrons in conduction and does not freeze them in the core. We have used a 161616 standard sampling of the Brillouin zone (BZ) and a plane waves (PW) cutoff of 65Β Ha for the DFT-PBE starting calculation, as well as a Gaussian smearing of 0.01 Hartree. In the subsequent calculations, the BZ sampling has been reduced to 666, whereas the PW cutoffs have been reduced to 40 Ha for the representation of the wave functions and for the exchange self-energy, and further down to 15 Ha for the correlation self-energy and the screening. We included 187 bands for the calculation of the screening and 225 for the self-energy (see Sect.Β F in the AppendixΒ for convergence studies). We have carried just only one iteration (i.e.Β ) on top of PBE, using a Godby-Needs plasmon-pole model [66] and a shift of 0.1Β eV to avoid poles/divergences. The final band plots, total and projected density of states (DOS and PDOS), as well as Fermi surfaces, were interpolated from the 666 to a denser -mesh by the Wannier90 code using a set of 67 projected Wannier functions (PWFs) with the closest possible character to atomic orbitals, namely O-, Ni-, La- and La-. The band orbital characters and PDOS identified by these PWFs are almost coincident with what one would obtain by projecting Bloch wave functions directly onto real atomic orbitals, which is not the case when selecting a set of maximally-localized Wannier functions (MLWFs) (see App.Β D). In the bandplots showing projected orbital character, the area of the dots scales as the orbital contribution squared, so that the width of the line is proportional to it.
All calculations have been done with a Gaussian smearing temperature of 0.01Β Ha to ease self-consistency convergence and optical absorption onset stability (which in any case has little relevance on main plasmons, energy-loss and screening ). For the most critical case at Β GPa we have checked that a calculation with enforced at Β Ha in both DFT and calculations provides the same physical picture, and in particular the same Fermi surface as in Fig.Β 1b, with only minor adjustments of less than Β meV on band positions, mostly due to the fact that the chemical potential at larger temperature is a less accurate estimate for the Fermi energy (see App.Β C).
@@ II.2 Crystal structure and Brillouin zone
To determine the crystal structure of La3Ni2O7 at ambient pressure, and then in the superconducting range at 14, 29.5 and 40 GPa, we have performed DFT PBE structural relaxations which provided the I4/mmm space group as the lowest energy structure at all considered pressures, except at 0Β GPa where the Amam structure is more stable. Nevertheless, our calculations have all referred to the I4/mmm structure even at ambient pressure to ease the comparison and better follow the evolution of the electronic structure with pressure, from the superconducting range down to 0Β GPa. For all the cases we have used the relaxed lattice parameters and the internal atomic position out of our PBE relaxations, except for the case at 29.5Β GPa which we treated separately.
@@
Indeed, 29.5Β GPa is the most studied case in the theoretical literature, so that we use it as a benchmark case to compare with previous calculations. Thus, for the sake of comparison, we decided to stick to the lattice parameters and atomic positions used in most of previous works [45, 24]. Therefore, we consider a tetragonal structure with Β Γ Β and Β Γ Β as Sakakibara et al.Β [24]. In fact, instead of the initial space-group symmetry proposed for La3Ni2O7 at 29.5Β GPa, our calculations predict the space-group symmetry with lattice parameters that are in good agreement with the experimental ones [67]. Further, in Ref.Β 45, it was noticed that the electronic structure of La3Ni2O7 at 29.5Β GPa is extremely sensitive to internal atomic positions and that the experimental ones reported by Sun et al.Β from the X-ray diffraction [1] likely yield artifacts at the Fermi level. We provide a detailed analysis of this question in AppendixΒ B.
@@
To simulate epitaxial strain, we enforced the tetragonal lattice parameter to the experimentally observed value of 3.77Β Γ Β [5] and let the system relax in the direction, which provided Β Γ . This crystal structure corresponds to a bulk non-hydrostatic planar stressed system, or equivalently a system pulled along the axis, that we will use to simulate an epitaxially strained thin film. All our crystal structures are detailed in App.Β A.
@@
All our electronic structures are reported in the body-centered tetragonal (BCT) Brillouin zone (BZ) which we report in Fig.Β 1c by black lines, while the blue line shows the path followed in our band plots with -point labels using the convention of Hinuma et al. Β [68]. Notice that in this convention, the X and M -points do not correspond to the canonical ones used in the 2D BZ of the square lattice widely used in the model literature: X is at the BZ corner in the nodal direction and M lies in the plane. We also indicate in grey the BZ corresponding to the conventional cell, namely simple tetragonal, which ease the comparison with 2D BZs. In the same Fig.Β 1c, we report also the simple orthorhombic BZ and associated -path to compare with the ARPES experiment [56] which referred to the tilted Amam structure needing a supercell compared to ours; and finally the face-centered orthorhombic BZ (and associated -path) to compare with the calculation of Christiansson et al.Β [45] who referred to the crystal structure.
@@ III Results
We start discussing the electronic structure of La3Ni2O7 at the pressure of 29.5Β GPa over which most of previous theoretical calculations have been performed. Our PBE and Fermi surfaces are shown in Fig.Β 1a and 1b, while Fig.Β 1d shows the corresponding band structures. In the same figure, selected orbital character contributions are plotted on the band structure, as well as the projected densities of states (PDOS) and the total density of states (DOS).
We briefly mention that, like it was already found [69, 7, 8] in infinite-layer nickelates, e.g.Β LaNiO2, correlations shift 2Β eV up the flat bands with predominant La-4 character, moving them further away from the Fermi level and out of the low-energy region, and shift 1Β eV down the O-2 manifold, leading to an increased ionic charge-transfer energy between Ni-3 and O-2 states. These aspects were previously detailed [69] and will not be again discussed here where we focus on the low-energy region.
@@ III.1.1 Ni Fermi surface sheets
The Fermi level is dominated by bands with Ni-3 character. The DFT PBE Fermi surface consists of three sheets, labeled , , and in previous literature [18]. The sheet is a large cylindrical sheet centered around ; is the sheet with the characteristic shape of the cuprate Fermi surface; and is a small hole pocket located at the Brillouin zone (BZ) corners which is highly dispersive along . The and sheets are primarily of character, though they show significant hybridization with orbitals in the anti-nodal direction, as well as with O-2 states (see Fig.Β 2b). In contrast, the pocket has an almost pure character and is consistently present in PBE (or LDA) calculations.
@@
The Fermi surface, in contrast, has no hole-pockets, as shown in Fig.Β 1b. This result is totally consistent with experimental ARPES data [56], with respect to which we provide a detailed comparison in Sec.Β III.2 below, since it was performed at ambient pressure. This difference between DFT and is due to a slight downward shift (7080 meV) of the lower Ni-3 band along the XP direction (see Fig.Β 1d). The absence of the pocket in experiment has been discussed also in the literature, and several semi-empirical corrections have been shown to reconcile theoretical calculations with the ARPES experiment. These methods either include the use of a weighted Fock exchange operator, as by HSE [54] or any other hybrid approach; or place more strength on correlation by introducing a Hubbard repulsion term of Β eV on Ni-3 electrons, like it was done in the calculation of Yang et al.Β [56], with the effect to lower this band. On the basis of this fact, Yang et al.Β [56] concluded that La3Ni2O7 is a strongly correlated system. Here we show that, a calculation using physical quasiparticle energies out of a many-body self-energy, even in a non self-consistent approximation, is enough to get rid of the sheet at the Fermi level and achieve good agreement with the experiment. Our results are in line with the findings of Ryee et al.Β [55] who excluded both a transition to a Mott insulator or bad metal scenario.
@@
The disappearance of the hole pocket has also been reported in a previous work [45] (see Fig.Β 1c for -point correspondence). They performed a calculation based on a reduced four-orbital tight-binding model restricted to the Ni- manifold. This model does not include the La-5 band, which, as we will discuss in the next section, plays a non-negligible role in the low energy spectrum. Their model also employs an effective screening that is renormalized to account for the missing degrees of freedom, and refers to the Fmmm structural phase.
@@
Despite these significant methodological differences, our results are in remarkable agreement with theirs regarding the absence of the pocket. However, Christiansson et al.Β do not further comment on this disappearance in their calculation, because their subsequent EDMFT simulationsβperformed on top of both DFT and βreintroduce and even enhance the spectral weight of the lower Ni-3 band at the Fermi level. On this basis, they conclude that βthe dominance of the orbital distinguishes La3Ni2O7 from the infinite-layer nickelates or the cuprates,β which are instead dominated by character. This interpretation is not supported by our results (see Sec.Β III.1.4 for the spectral weights), which show a predominantly character at low energy. Our findings rather align with the recent calculations of You et al.Β [15].
@@
Interestingly, Ryee et al.Β [55] found the pocket in both DFT and dynamical mean-field theory (DMFT) calculations, but not in cluster-DMFT (CDMFT) which is therefore more in agreement with the finding. This is an indication that non-local correlations, as provided by a CDMFT or a -dependent self-energy, are important to describe such modification of the Fermi surface topology. However, CDMFT also find an important modification to the cuprate-shape sheet which loses most of its 1D character (see below). On this point is more in agreement with DFT and local DMFT rather than with CDMFT.

### v2

@@ Abstract
The recent discovery of superconductivity in La3Ni2O7 at a critical temperature above 80 K points to a nonconventional pairing mechanism in nickelates as in cuprates, possibly due to electronic correlations. We have calculated from first principles the electronic structure of La3Ni2O7 under the effect of pressure and epitaxial strain including correlations by the approximation to the many-body self-energy. We find that the Fermi surface is composed of a characteristic cuprate-shape sheet plus a nickelate-specific cylinder , both from Ni orbitals, with a nonnegligible drop in the quasiparticle weight and an effective 1D character. This topology results from a delicate balance between the Ni-3 hole pocket , which is suppressed by correlations, and an emerging La-5 electron pocket induced by both correlation and pressure/strain effects and whose role at low energy has been neglected so far. Unlike cuprates, the electronic structure of La3Ni2O7 is already correctly described from ab initio and in agreement with the experiment without the need to introduce Hubbard adjustable parameters or to invoke a strongly correlated physics.
@@ I Introduction
The recent discovery of superconductivity (SC) in Ruddlesden-Popper (RP) nickelates [1, 2, 3] has introduced a new direction in the field of high-temperature superconductivity. Notably, superconductivity has been observed in bulk samples of the bilayer La3Ni2O7 under high pressure [1], with the superconducting phase forming a right-triangular region spanning 14 to 80 GPa, and reaching a maximum onset critical temperature of approximately 80 K at 18 GPa [4]. More recently, superconductivity has also been reported at K in La3Ni2O7 thin films under epitaxial compressive strain at ambient pressure [5, 6]. While the superconducting mechanism is still debated in infinite-layer LnNiO2 nickelates [7, 8, 9, 10, 11], where critical temperatures do not overstep 30 K [12, 13], the high critical temperature observed in La3Ni2O7 is difficult to reconcile with a BCS electron-phonon pairing mechanism [14, 15], thus pointing to unconventional superconductivity like in cuprates. Indeed, cuprates and nickelates share many analogies: they are both organized in a layered crystal structure with CuO2 or NiO2 planes; and present the same “cuprate-shape” Fermi surface sheet of main 3 character, although nickelates bilayer and trilayer show a strong Ni-3 hybridization in the antinodal direction. However, there are also important differences: the constant presence of an antiferromagnetic insulating parent phase in cuprates has no analogous in nickelates. For what concerns the electronic structure, the fact that in nickelates the Ni-3 states are closer to the Fermi level than in cuprates, gives rise to at least another Fermi sheet and possibly to a multiband character superconductivity. Therefore, although both unconventional, the pairing mechanism could be possibly different in cuprates and nickelates.
@@
In this work we study from ab initio the electronic structure of La3Ni2O7 including correlations at the level of the approximation to the self-energy [59, 60, 61, 62]. In contrast to the Kohn-Sham eigenvalues used in density-functional theory (DFT), which are the energies of a fictitious noninteracting system, quasiparticle energies have a direct physical interpretation as the excitation energies measured in ARPES. The approximation takes into account electronic exchange and correlations without introducing any adjustable parameter, neither in the DFT functional (e.g. an hybrid mixing parameter) nor in the Hamiltonian (e.g. a Hubbard interaction term). The weight of correlation effects and their importance in this system is revealed by a comparison between the electronic structure and the DFT band plot. However, in this work we applied the one-iteration approach which can keep some reminiscence of the DFT starting point.
We also investigate the electronic structure evolution of bulk La3Ni2O7 with respect to pressure, as well as in presence of in-plane constraints that simulate a thin film grown on a substrate. Our purpose is to highlight aspects which might be relevant to explain superconductivity that in this system is induced by pressure without doping. The focus is at the Fermi level but also at higher energy features which are less affected by the estimated 0.1 eV best absolute accuracy, and which might be relevant in the competition between different pairing symmetries according to some proposed superconducting mechanisms [24]. Our work describes some of the electronic structure features expected to be relevant by theoretical models so far proposed [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48], and provides also indications for future works.
The paper is organized as follows: after introducing our methods in Sec. II, we will first compare our calculations with previous ab initio works done at 29.5 GPa with the crystal structure used by almost all DFT as well as two calculations [45, 15] (Sec. III.1). We will then show the evolution with pressure of the electronic structure for the bulk crystal in Sec. III.2, and finally introduce in Sec. III.3 our results for the structure with epitaxial strain of 1.8%. Our results will be compared with the available experimental data from photoemission (ARPES) experiments.
@@ II.1 Numerical methods
Our methodology relies as a first step on density-functional theory (DFT) calculations in the PBE approximation [63], using the plane waves code Abinit [64]. We use norm-conserving pseudopotentials from the PseudoDojo table library [65]. The La pseudopotential takes 4 electrons in conduction and does not freeze them in the core. We have used a 161616 standard sampling of the Brillouin zone (BZ) and a plane waves (PW) cutoff of 65 Ha for the DFT-PBE starting calculation, as well as a Gaussian smearing of 0.01 Hartree. In the subsequent calculations, the BZ sampling has been reduced to 666, whereas the PW cutoffs have been reduced to 40 Ha for the representation of the wave functions and for the exchange self-energy, and further down to 15 Ha for the correlation self-energy and the screening. We included 187 bands for the calculation of the screening and 225 for the self-energy (see Sect. G in the Appendix for convergence studies). We have carried just only one iteration (i.e. ) on top of PBE, using a Godby-Needs plasmon-pole model [66] and a shift of 0.1 eV to avoid poles/divergences. The final band plots, total and projected density of states (DOS and PDOS), as well as Fermi surfaces, were interpolated from the 666 to a denser -mesh by the Wannier90 code using a set of 67 projected Wannier functions (PWFs) with the closest possible character to atomic orbitals, namely O-, Ni-, La- and La-. The band orbital characters and PDOS identified by these PWFs are almost coincident with what one would obtain by projecting Bloch wave functions directly onto real (pseudo) atomic orbitals, as we have checked using the Quantum Espresso code with equivalent parameters, which is not the case when selecting a set of maximally localized Wannier functions (MLWFs) (see App. E). In the bandplots showing projected orbital character, the width of the dots is proportional to the orbital contribution.
All calculations have been done with a Gaussian smearing temperature of 0.01 Ha to ease self-consistency convergence and optical absorption onset stability (which in any case has little relevance on main plasmons, energy-loss and screening ). For the most critical case at GPa we have checked that a calculation with enforced at Ha in both DFT and calculations provides the same physical picture, and in particular the same Fermi surface as in Fig. 1b, with only minor adjustments of less than meV on band positions, mostly due to the fact that the chemical potential at larger temperature is a less accurate estimate for the Fermi energy (see App. D).
@@ II.2 Crystal structure and Brillouin zone
To determine the crystal structure of La3Ni2O7 at ambient pressure, and then in the superconducting range at 14, 29.5 and 40 GPa, we have performed DFT PBE structural relaxations which provided the I4/mmm space group as the lowest energy structure at all considered pressures, except at 0 GPa where the Amam structure is more stable. Nevertheless, our calculations have all referred to the I4/mmm structure even at ambient pressure to ease the comparison and better follow the evolution of the electronic structure with pressure, from the superconducting range down to 0 GPa. For all the cases we have used the relaxed lattice parameters and the internal atomic position out of our PBE relaxations, except for the case at 29.5 GPa which we treated separately.
@@
Indeed, 29.5 GPa is the most studied case in the theoretical literature, so that we use it as a benchmark case to compare with previous calculations. Thus, for the sake of comparison, we decided to stick to the lattice parameters and atomic positions used in most of previous works [45, 24]. Therefore, we consider a tetragonal structure with Å and Å as Sakakibara et al. [24]. In fact, instead of the initial space-group symmetry proposed for La3Ni2O7 at 29.5 GPa, our calculations predict the space-group symmetry with lattice parameters that are in good agreement with the experimental ones [67]. Further, in Ref. 45, it was noticed that the electronic structure of La3Ni2O7 at 29.5 GPa is extremely sensitive to internal atomic positions and that the experimental ones reported by Sun et al. from the x-ray diffraction [1] likely yield artifacts at the Fermi level. We provide a detailed analysis of this question in App. B.
@@
To simulate epitaxial strain, we enforced the tetragonal lattice parameter to the experimentally observed value of 3.77 Å [5] and let the system relax in the direction, which provided Å. This crystal structure corresponds to a bulk nonhydrostatic planar stressed system, or equivalently a system pulled along the axis, that we will use to simulate an epitaxially strained thin film. All our crystal structures are detailed in App. A.
@@
All our electronic structures are reported in the body-centered tetragonal (BCT) Brillouin zone (BZ) which we report in Fig. 1c by black lines, while the blue line shows the path followed in our band plots with -point labels using the convention of Hinuma et al. [68]. Notice that in this convention, the X and M -points do not correspond to the canonical ones used in the 2D BZ of the square lattice widely used in the model literature: X is at the BZ corner in the nodal direction and M lies in the plane. We also indicate in grey the BZ corresponding to the conventional cell, namely simple tetragonal, which ease the comparison with 2D BZs. In the same Fig. 1c, we report also the simple orthorhombic BZ and associated -path to compare with the ARPES experiment [56] which referred to the tilted Amam structure needing a supercell compared to ours; and finally the face-centered orthorhombic BZ (and associated -path) to compare with the calculation of Christiansson et al. [45] who referred to the crystal structure.
@@ III Results
We start discussing the electronic structure of La3Ni2O7 at the pressure of 29.5 GPa over which most of previous theoretical calculations have been performed. Our PBE and Fermi surfaces are shown in Fig. 1a and 1b, while Fig. 1d shows the corresponding band structures. In the same figure, selected orbital character contributions are plotted on the band structure, as well as the projected densities of states (PDOS) and the total density of states (DOS).
We briefly mention that, like it was already found [69, 7, 8] in infinite-layer nickelates, e.g. LaNiO2, correlations shift 2 eV up the flat bands with predominant La-4 character, moving them further away from the Fermi level and out of the low-energy region, and shift 1 eV down the O-2 manifold, leading to an increased ionic charge-transfer energy between Ni-3 and O-2 states. These aspects were previously detailed [69] and will not be again discussed here where we focus on the low-energy region.
@@ III.1.1 Ni Fermi surface sheets
The Fermi level is dominated by bands with Ni-3 character. The DFT PBE Fermi surface consists of three sheets, labeled , , and in previous literature [18]. The sheet is a large cylindrical sheet centered around ; is the sheet with the characteristic shape of the cuprate Fermi surface; and is a small hole pocket located at the Brillouin zone (BZ) corners which is highly dispersive along . The and sheets are primarily of character, though they show significant hybridization with orbitals in the antinodal direction, as well as with O-2 states (see Fig. 2b). In contrast, the pocket has an almost pure character and is consistently present in PBE (or LDA) calculations.
@@
The Fermi surface, in contrast, has no hole-pockets, as shown in Fig. 1b. This result is totally consistent with experimental ARPES data [56], with respect to which we provide a detailed comparison in Sec. III.2 below, since it was performed at ambient pressure. This difference between DFT and is due to a slight downward shift (7080 meV) of the lower Ni-3 band along the XP direction (see Fig. 1d). The absence of the pocket in experiment has been discussed also in the literature, and several semi-empirical corrections have been shown to reconcile theoretical calculations with the ARPES experiment. These methods either include the use of a weighted Fock exchange operator, as by HSE [54] or any other hybrid approach; or place more strength on correlation by introducing a Hubbard repulsion term of eV on Ni-3 electrons, like it was done in the calculation of Yang et al. [56], with the effect to lower this band. On the basis of this fact, Yang et al. [56] concluded that La3Ni2O7 is a strongly correlated system. Here we show that, a calculation using physical quasiparticle energies out of a many-body self-energy, even in a non self-consistent approximation, is enough to get rid of the sheet at the Fermi level and achieve good agreement with the experiment. Our results are in line with the findings of Ryee et al. [55] who excluded both a transition to a Mott insulator or bad metal scenario.
@@
The disappearance of the hole pocket has also been reported in a previous work [45] (see Fig. 1c for -point correspondence). They performed a calculation based on a reduced four-orbital tight-binding model restricted to the Ni- manifold. This model does not include the La-5 band, which, as we will discuss in the next section, plays a nonnegligible role in the low energy spectrum. Their model also employs an effective screening that is renormalized to account for the missing degrees of freedom, and refers to the Fmmm structural phase.
@@
Despite these significant methodological differences, our results are in remarkable agreement with theirs regarding the absence of the pocket. However, Christiansson et al. do not further comment on this disappearance in their calculation, because their subsequent EDMFT simulations—performed on top of both DFT and —reintroduce and even enhance the spectral weight of the lower Ni-3 band at the Fermi level. On this basis, they conclude that “the dominance of the orbital distinguishes La3Ni2O7 from the infinite-layer nickelates or the cuprates,” which are instead dominated by character. This interpretation is not supported by our results (see Sec. III.1.4 for the spectral weights), which show a predominantly character at low energy. Our findings rather align with the recent calculations of You et al. [15].
@@
Interestingly, Ryee et al. [55] found the pocket in both DFT and dynamical mean-field theory (DMFT) calculations, but not in cluster-DMFT (CDMFT) which is therefore more in agreement with the finding. This is an indication that nonlocal correlations, as provided by a CDMFT or a -dependent self-energy, are important to describe such modification of the Fermi surface topology. However, CDMFT also find an important modification to the cuprate-shape sheet which loses most of its 1D character (see below). On this point is more in agreement with DFT and local DMFT rather than with CDMFT.